# Investigating cannibalistic millisecond pulsar binaries using MESA : new constraints from pulsar spin and mass evolution

### v1

@@ 2.3.1 Accretion regime, $r_{\rm mag}<(r_{\rm cor},r_{\rm lc})$
If the mass-transfer rate is high enough that the magnetospheric radius is below the co-rotation radius (following Equation 8), we assume that accretion can proceed unimpeded. Since the magnetic field is weak, we assume that the inner accretion disc extends down to the NS surface. Hence, in Equation 6, $R_{\rm NS}$ is equal to the NS radius, which we take as 12.5 km (for example, Raaijmakers et al., 2021). Since, the increase in angular momentum comes from accreted material, we take $\dot{M}_{\rm NS}=\epsilon\dot{M}_{\rm c}=(1-\beta)\dot{M}_{\rm c}$, $\epsilon$ being the accretion efficiency. We also assume a $\kappa$ of unity, similar to Chattopadhyay et al. (2020).
@@ 2.3.3 Ejection regime and pulsar wind phase
If the magnetic field is strong enough, or the mass-accretion rate is low enough, the magnetosphere extends far outside of the light cylinder radius. Any accretion or transfer of angular momentum is prevented and the pulsar spin (and the corresponding pulsar magnetic field) is evolved as that for an isolated pulsar (see Section 2.2). We use two criteria to determine when the ejection regime and the consequent pulsar wind phase begins. First, the pulsar has accreted enough matter to spin-up to MSP-like spins (we use 30 ms as the identifying upper limit). The second condition for the onset of the pulsar wind phase is that the mass-transfer rate has decreased below the criteria for thermal-viscous instabilities and can no longer sustain a persistent accretion disc (see Equation 3), under the effect of X-ray irradiation during the RLO phase.
Once the ejection phase begins and as the pulsar spins down, the pulsar emits an energetic wind with a certain luminosity, known as the spin-down luminosity, which is defined using the spin properties of the pulsar as follows,
@@
In the above equation, the initial values of the NS spin ($P_{\rm spin}$) and change in NS spin with time ($\dot{P}_{\rm spin}$) are determined by the preceding accretion phase (see Section 2.3.1) and follow the spin evolution of a detached pulsar (see Section 2.2). The irradiation from the pulsar wind drives a mass loss from the companion surface ($\dot{M}_{\rm c}$), defined by Stevens et al. (1992),
$\dot{M}_{\rm c}=-f_{\rm pulsar}\times\frac{L_{\rm pulsar}}{2v^{2}_{\rm esc}}\bigg{(}\frac{R_{\rm c}}{a}\bigg{)}^{2},$ (10)
where $f_{\rm pulsar}$ is the efficiency that describes the transfer of energy from the incoming spin-down luminosity to the generated wind loss. It also stands for any geometrical beaming effects that might be present during the process. The other parameters in the above equation include $L_{\rm pulsar}$ as the spin-down luminosity, $v_{\rm esc}$ as the escape velocity from the companion star surface, $R_{\rm c}$ as the companion radius, and $a$ as the binary separation.
@@ 2.4 Observations used for comparisons to simulations
To compare our simulations to observations we use a catalogue of BWs, RBs, and huntsmen (including candidates). A total of 57 spiders are used, 52 of which are catalogued in Linares & Kachelrieß (2021, and references therein). The more recent ones are from various other works in the literature, and an updated catalogue has been prepared by Nedreaas (2024, and references therein). The reported values for the companion masses are in terms of the minimum mass (denoted as $M_{\rm c,min}$), calculated assuming a pulsar mass of $1.4\,{\rm M}_{\odot}$ and inclination of $90\degree$ from observations of radio pulses. Only for one of the huntsmen source in the catalogue have radio pulses been confirmed and $M_{\rm c,min}$ estimated (1FGL J1417.7–440 Camilo et al., 2016). While the other huntsmen, 2FGL J0846.0+2820, only has Fermi Large Area Telescope (LAT) observations (Swihart et al., 2017).
@@ 3 Results
Figure 1: Orbital period evolution versus companion mass of the simulated binaries, for $\beta=0.3$, having an initial NS of mass $1.3\,{\rm M}_{\odot}$ and $f_{\rm pulsar}$ values of 0.0 (top-left), 0.05 (top-right), 0.1 (bottom left), and 0.5 (bottom-right). $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. The colorbar shows the changing surface abundance of the companion. We also compare our tracks to estimated $M_{\rm c,min}$ and observed $P_{\rm orb}$ for BWs (black crosses) and RBs (red crosses), and huntsmen spiders (green crosses), including whether H-lines have been detected (filled circles) or not (filled triangles) or if it is unknown (crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017) (surface H is unknown). The three tidarrens (J1311-3430, J1653-0159, and J0636+5128) are also shown. The initial binaries have a MS star (range of 1.0 to 4.5 ${\rm M}_{\odot}$) and initial periods of 0.6 to 4.0 d.
Figure 2: Distribution of the final NS spins ($P^{f}_{\rm spin}$) in converging binaries, for $\beta=0.0$ (top), $\beta=0.3$ (middle), and $\beta=0.7$ (bottom). We group the binaries by the initial NS mass ($M^{i}_{\rm NS}$), and if the final mass was above $2.5\,{\rm M}_{\odot}$ (hatched histograms, with faded colours) or below it (step histograms, with brighter colours). We also compare to the observed spins of BWs (shown by the gray shaded area).
@@ 3.1 Mass-accretion efficiency
We investigate the formation of different branches of compact MSP binaries using detailed evolutionary sequences. Figures 1 and 13 show the evolutionary tracks with $\beta$ of 0.3 and 0.7, respectively, each with different values of $f_{\rm pulsar}$. The tracks correspond to initial NS masses of 1.3 M and have final NS spins  $\lesssim\rm\,30\,ms$. While $\beta=0.7$ reproduces the very compact spiders, it cannot explain a large part of the binary parameter space where these systems are observed. This behavior is irrespective of the pulsar wind efficiencies. Similar behavior is seen for huntsmen spiders ($P_{\rm orb}$ 1.0 d), as $\beta=0.7$ does not reproduce the small population of observed huntsmen spiders. With $\beta$ of 0.3, a large part of the observed parameter space is explained and on increasing $f_{\rm pulsar}$, more and more BWs and RBs’ observed orbital properties are reproduced. The reason for the difference between $\beta=0.3$ and $\beta=0.7$ can be understood by looking at Equation 10 which shows that the mass lost due to pulsar wind irradiation is directly proportional to the pulsar spin-down luminosity, which in turn depends on the pulsar spin at the end of the RLO phase (Equation 9). Since the preceding accretion phase dictates the spin-up of the pulsar, we can compare the resulting pulsar spins in our grids with observations to see why more accretion helps.
@@
We find that all the MESA  tracks that pass through the RB region of the observed parameter space, end up in the BW region (see Figure 1). The only exception is for one track corresponding to $f_{\rm pulsar}=0.5$, that ends with $M_{\rm c}$ of around 0.15 ${\rm M}_{\odot}$ (discussed below in Section 3.4). Hence, to study the effect of the assumed $\beta$ on pulsar spin-up, we compare only the observed BWs to the final pulsar spins from the grid in Figure 2, corresponding to $f_{\rm pulsar}=0.0$ and $\beta=$ 0.0, 0.3, and 0.7. The figure shows the different binaries with different initial NS masses, and also separates out the cases where the final NS mass exceeds 2.5 M, which is an approximate upper limit for the NS mass. NSs with final masses above 2.5 M are considered to have collapsed into BHs and would no longer be observed as radio pulsars. The value $\beta=0.0$ leads to a majority of the NSs gaining spins in the sub-MSP range ($\lesssim 1$ ms) and having final NS masses $\gtrsim 2.5\,{\rm M}_{\odot}$.
@@
With a $\beta$ of $0.7$ (or an accretion efficiency of $0.3$), the minimum spin period that is reached in all the grids is 1.14 ms, consistent with the fastest spinning Galactic BW observed so far (PSR J0952–0607, with a spin period of 1.41 ms; Bassa et al., 2017). However, these short spins correspond to the case where the final NS mass exceeds the maximum NS mass limit. Removing these, the minimum spin period for $\beta=0.7$ is 3.2 ms, which is more than twice the minimum observed BW spin. Looking at the case of $\beta=0.3$ in Figure 2, all the observed spins can be reproduced. Comparing the minimum simulated spin period for $\beta=0.3$ (1.38 ms) and the corresponding final pulsar mass ($2.36\,{\rm M}_{\odot}$) with PSR J0952–0607, we find that our simulations can reproduce well this extreme BW pulsar (its estimated pulsar mass is $2.35^{+0.17}_{-0.17}\,{\rm M}_{\odot}$; Romani et al., 2022). There are again some binary tracks that ended with final spins below 1 ms, in the sub-MSP regime (all with $M^{i}_{\rm NS}>1.3\,{\rm M}_{\odot}$). The NSs with $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$ gained a maximum of $1.06\,{\rm M}_{\odot}$ and did not spin-up to sub-MSP regime (final spin was 1.37 ms). All the binaries where the final NS mass exceeded the maximum NS mass limit, were spun up to sub-milliseconds. Hence, the mass required to spin-up a NS to sub-millisecond range is high enough to collapse it into a BH. Additionally, the sub-millisecond regime is mainly dominated by initially more massive NSs, 1.8 and 2.0 M (see Figure 2), while the initially less massive NSs are not always able to accrete enough matter to enter the sub-millisecond regime. This is because more massive NSs leads to more a stable and longer lasting mass transfer phase (Misra et al., 2020). Comparing the final spin distributions between $\beta=0.3$ and $\beta=0.7$, we can see that $\beta=0.7$ results in slower spinning NS, which would then have lower spin-down luminosities as $L_{\rm pulsar}\propto P^{-3}_{\rm spin}$, and not affecting the orbital evolution as much. Increased wind mass loss from the companion would lead to increased orbital widening during the radio pulsar phase.
@@
Figure 3: Evolution of orbital period versus companion mass for selected binaries from our simulated grids ($\beta=0.3$, $f_{\rm pular}=0.0$, and $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$). $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. The three different kinds of binaries that we discuss in the text are studied here, case A ($P^{i}_{\rm orb}=1.0\rm\,d$), near-$P_{\rm bif}$ ($P^{i}_{\rm orb}=2.4\rm\,d$), and diverging ($P^{i}_{\rm orb}=2.6\rm\,d$). Also shown is the binary that goes down to the orbital period of 19 min (orange curve) for comparison, which is the minimum in the grid with $f_{\rm pulsar}=0.0$. We show the start of the first RLO phase as circles and the formation of a fully convective star as squares (which happens only for two of the tracks shown). We also compare our tracks to estimated $M_{\rm c,min}$ and observed $P_{\rm orb}$ for BWs (faint black crosses), RBs (faint red crosses), and huntsmen spiders (faint green crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017). The tidarrens, namely, J1311–3430, J1653–0159, and J0636+5128, are highlighted as pink, blue, and black crosses, respectively.
@@ 3.2 Orbital evolution and surface hydrogen abundance
The top left panel in Figure 1 shows the evolutionary tracks with no pulsar wind irradiation ($f_{\rm pulsar}=0.0$). The colorbar corresponds to the surface H abundance of the companion, which can be compared to optical spectroscopic observations of spiders. We mark in Figure 1 spiders where H lines have not been detected in their optical spectra (filled triangles) and those where H lines are detected (filled circles). Two of the most compact BWs, PSR J1311–3430 (Pletsch et al., 2012; Romani et al., 2012) and PSR J1653–0159 (Kong et al., 2014; Romani et al., 2014), with negligible surface H abundance, are easily reproduced by our evolutionary tracks. These sources also form a sub-category of BWs, called tidarrens, with $P_{\rm orb}\lesssim 2$ hr and $M_{\rm c}\lesssim 0.015\,{\rm M}_{\odot}$ (Romani et al., 2016). They are reproduced regardless of the value of $f_{\rm pulsar}$. Their evolutionary tracks also pass through the ultra-compact X-ray binary region (UCXB), with orbital periods less than 1 hr, with the minimum orbital period being 30.8 min. Most UCXBs are also considered to be H-deficient due to their small orbits (Nelson et al., 1986; Nelemans et al., 2006), similar to our binary evolution tracks for tidarrens. Hence, tidarrens (the most compact BWs with the lightest companions) passed through a UCXB phase in their previous evolution. Additionally, we also see these systems reproduced in Figure 13 (with $\beta=0.7$), since not a lot of accretion is needed to form these sources.
@@
We can understand their evolution better in Figure 3, where we show tracks that pass through the tidarren region (solid gray line, with an $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=$2.4 d) and compare these with other tracks with initially narrower and wider orbits. All the tracks shown in the figure have an $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$. We also show a binary that goes well into the UCXB region during its evolution, to reach the overall minimum orbital period (19 min) in the grid slice (with an $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=$2.6 d). With increasing orbital periods, the binaries go from starting RLO with the companion on the MS (at $P_{\rm orb}=0.5$ d, also called case A RLO), to starting RLO when the companion has exhausted core-H (also called case B RLO, at $P_{\rm orb}=1.1$ d). This transition occurs at the bifurcation period ($P_{\rm bif}$), because companions in wider orbits have more time to evolve and move out of the MS. In the presented example, $P_{\rm bif}$ is between $P_{\rm orb}=0.93$ d (with $P^{i}_{\rm orb}=2.4$ d and $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$) and $P_{\rm orb}=1.1$ d (with $P^{i}_{\rm orb}=2.6$ d and $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$). We term systems just below $P_{\rm bif}$ as near-$P_{\rm bif}$ binaries, since similar studies have shown that the formation of UCXBs from LMXBs happens just below $P_{\rm bif}$ (Podsiadlowski et al., 2002, 2003; Gossage et al., 2023).
@@
For binary evolution governed only by conservative mass transfer, the orbit would expand as long as the mass is being transferred from the less massive to the more massive star, while it would contract if the companion star is more massive. With the involvement of magnetic braking, this evolution is not as simple. In NS binaries with MS companions having masses around $1\,{\rm M}_{\odot}$, orbital angular momentum evolution is dominated by magnetic braking which contracts the orbits. Since case A RLO (corresponding to the black curve in Figure 3) proceeds on a nuclear timescale during the evolution of the stellar MS, any mass loss via winds is not fast enough to dominate over magnetic braking. Additionally, in compact binaries ($\lesssim 1\rm\,d$), angular momentum losses due to gravitational wave radiation are also present and further contract the orbit. Once the companion in case A binaries decreases to about 0.2 to 0.3 M, it develops a fully convective structure and magnetic braking stops operating (Rappaport et al., 1983; Spruit & Ritter, 1983; Pylyser & Savonije, 1989; Podsiadlowski et al., 2002). This is most prominent for the binary in Figure 3 with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=1$ d; there is a decrease in the orbit without any change in mass in the companion mass. The companion detaches from its Roche lobe as the main process driving RLO (the first RLO phase for this binary) has stopped operating, until it fills its Roche lobe again due to gravitational wave radiation continuing to contract the orbit (which is the drop in orbital period seen in the figure). The gravitational wave radiation sustains this second RLO phase and the orbit eventually expands.

### v2

@@ 2.3.1 Accretion regime, $r_{\rm mag}<(r_{\rm cor},r_{\rm lc})$
If the mass-transfer rate is high enough that the magnetospheric radius is below the co-rotation radius (following Equation 8), we assume that accretion can proceed unimpeded. Since the magnetic field is weak, we assume that the inner accretion disk extends down to the NS surface. Hence, in Equation 6, $R_{\rm NS}$ is equal to the NS radius, which we take as 12.5 km (for example, Raaijmakers et al., 2021). Since, the increase in angular momentum comes from accreted material, we take $\dot{M}_{\rm NS}=\epsilon\dot{M}_{\rm c}=(1-\beta)\dot{M}_{\rm c}$, $\epsilon$ being the accretion efficiency. We also assume a $\kappa$ of unity, similar to Chattopadhyay et al. (2020).
@@ 2.3.3 Ejection regime and pulsar wind phase
If the magnetic field is strong enough or the mass-accretion rate is low enough, the magnetosphere extends far outside of the light cylinder radius. Any accretion or transfer of angular momentum is prevented and the pulsar spin (and the corresponding pulsar magnetic field) is evolved as that of an isolated pulsar (see Section 2.2). We use two criteria to determine when the ejection regime and the consequent pulsar wind phase begins. First, the pulsar has accreted enough matter to spin up to MSP-like spins (we use 30 ms as the identifying upper limit). The second condition for the onset of the pulsar wind phase is that the mass-transfer rate has decreased below the criteria for thermal-viscous instabilities and can no longer sustain a persistent accretion disk (see Equation 3), under the effect of X-ray irradiation during the RLO phase.
Once the ejection phase begins and as the pulsar spins down, the pulsar emits an energetic wind with a certain luminosity, known as the spin-down luminosity, which is defined using the spin properties of the pulsar as follows:
@@
As this energetic wind hits the companion, it ablates the outer surface of the star, which, when observed, signifies the presence of a spider pulsar (Kluzniak et al., 1988; Ruderman et al., 1989b, a; Stappers et al., 1996b). In the above equation, the initial values of the NS spin ($P_{\rm spin}$) and change in NS spin with time ($\dot{P}_{\rm spin}$) are determined by the preceding accretion phase (see Section 2.3.1) and follow the spin evolution of a detached pulsar (see Section 2.2). The irradiation from the pulsar wind drives a mass loss from the companion surface ($\dot{M}_{\rm c,irr}$) defined by Stevens et al. (1992):
$\dot{M}_{\rm c,irr}=-f_{\rm pulsar}\times\frac{L_{\rm pulsar}}{2v^{2}_{\rm esc}}\bigg{(}\frac{R_{\rm c}}{a}\bigg{)}^{2},$ (10)
where $f_{\rm pulsar}$ is the efficiency that describes the transfer of energy from the incoming spin-down luminosity to the generated wind loss. It also stands for any geometrical beaming effects that might be present during the process. The other parameters in the above equation include $L_{\rm pulsar}$ as the spin-down luminosity, $v_{\rm esc}$ as the escape velocity from the companion star surface, $R_{\rm c}$ as the companion radius, and $a$ as the binary separation. The effect of this increased mass loss from the companion is described by the fast mode or Jean’s mode of mass transfer (Jeans, 1924; Soberman et al., 1997). This mode of mass transfer assumes that a small fraction of the stellar mass is lost in the form of fast, spherically symmetric winds that leave the binary without interacting with any binary components, escaping with the specific angular momentum of the mass-losing star. The orbital angular momentum ($J_{\rm orb}$) is defined as
$J_{\rm orb}=\mu\sqrt{Ga(M_{\rm c}+M_{\rm NS})},$ (11)
where $\mu$ is the reduced mass of the binary. The change in orbital angular momentum per unit of reduced mass ($J/\mu$) is negligible and the orbit expands accommodate the change in total mass (Huang, 1963). As the effect of irradiation will not necessarily be spherically symmetric, we have the efficiency term $f_{\rm pulsar}$ as described above (Stevens et al., 1992; Chen et al., 2013).
@@ 2.4 Observations used for comparisons to simulations
To compare our simulations to observations we use a catalog of BWs, RBs, and huntsmen (including candidates). A total of 57 spiders are used, 52 of which are cataloged in Linares & Kachelrieß (2021, and references therein). The more recent ones are from various other works in the literature, and an updated catalog has been prepared by Nedreaas (2024, and references therein). The reported values for the companion masses are in terms of the minimum mass (denoted as $M_{\rm c,min}$), calculated assuming a pulsar mass of $1.4\,{\rm M}_{\odot}$ and inclination of $90\degree$ from observations of radio pulses. Only for one of the huntsmen source in the catalog have radio pulses been confirmed and $M_{\rm c,min}$ estimated (1FGL J1417.7–440 Camilo et al., 2016). While the other huntsmen, 2FGL J0846.0+2820, only has Fermi Large Area Telescope (LAT) observations (Swihart et al., 2017).
@@ 3 Results
Figure 1: Orbital period evolution versus companion mass of the simulated binaries for $\beta=0.3$ with an initial NS of mass $1.3\,{\rm M}_{\odot}$ and $f_{\rm pulsar}$ values of 0.0 (top left), 0.05 (top right), 0.1 (bottom left), and 0.5 (bottom right). The term $M_{\rm c}$ is the mass of the companion to the NS, and $P_{\rm orb}$ is the orbital period. The color bar shows the changing surface abundance of the companion. We also compare our tracks to estimated $M_{\rm c,min}$ and observed $P_{\rm orb}$ for BWs (black crosses) and RBs (red crosses), and huntsmen spiders (green crosses), including whether H-lines have been detected (filled circles) or not (filled triangles) or if it is unknown (crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017) (surface H is unknown). The three tidarrens (J1311-3430, J1653-0159, and J0636+5128) are also shown. The initial binaries have a MS star (range of 1.0 to 4.5 ${\rm M}_{\odot}$) and initial periods of 0.6 to 4.0 d.
Figure 2: Evolution of the NS spins ($P_{\rm spin}$) versus the orbital periods ($P_{\rm orb}$) in converging binaries for $\beta=0.0$ (top), $\beta=0.3$ (middle), and $\beta=0.7$ (bottom). We distinguish the binaries if the final NS mass was above $2.5\,{\rm M}_{\odot}$ (faint gray curves) or below it, which are further grouped by two initial NS masses ($M^{i}_{\rm NS}$) 1.3 ${\rm M}_{\odot}$(blue curves) and 2.0 ${\rm M}_{\odot}$(orange curves). We also compare to the observed spins of BWs and RBs, shown by the black and red circles, respectively.
@@ 3.1 Mass-accretion efficiency
We investigate the formation of different branches of compact MSP binaries using detailed evolutionary sequences. Figures 1 and 13 show the evolutionary tracks with $\beta$ of 0.3 and 0.7, respectively, each with different values of $f_{\rm pulsar}$. The tracks correspond to initial NS masses of 1.3 ${\rm M}_{\odot}$and have final NS spins  $\lesssim\rm\,30\,ms$. While $\beta=0.7$ reproduces the very compact spiders, it cannot explain a large part of the binary parameter space where these systems are observed. This behavior is irrespective of the pulsar wind efficiencies. Similar behavior is seen for huntsmen spiders ($P_{\rm orb}$ 1.0 d), as $\beta=0.7$ does not reproduce the small population of observed huntsmen spiders. With $\beta$ of 0.3, a large part of the observed parameter space is explained and on increasing $f_{\rm pulsar}$, more and more BWs and RBs’ observed orbital properties are reproduced. The reason for the difference between $\beta=0.3$ and $\beta=0.7$ can be understood by looking at Equation 10, which shows that the mass lost due to pulsar wind irradiation is directly proportional to the pulsar spin-down luminosity and in turn depends on the pulsar spin at the end of the RLO phase (Equation 9). Since the preceding accretion phase dictates the spin-up of the pulsar, we can compare the resulting pulsar spins in our grids with observations to see why more accretion helps.
@@
We compared the observed BWs and RBs to the pulsar spin evolution from the grid in Figure 2, corresponding to $f_{\rm pulsar}=0.0$ and $\beta=$ 0.0, 0.3, and 0.7. The figure shows the binaries for two initial NS masses 1.3 ${\rm M}_{\odot}$and 2.0 ${\rm M}_{\odot}$, and also separates out the cases where the final NS mass exceeds 2.5 ${\rm M}_{\odot}$, which is an approximate upper limit for the NS mass. The NSs with final masses above 2.5 ${\rm M}_{\odot}$are considered to have collapsed into black holes (BHs) and would no longer be observed as radio pulsars. The value $\beta=0.0$ leads to a majority of the NSs with $M^{i}_{\rm NS}>1.3\,{\rm M}_{\odot}$ gaining spins in the sub-MSP range ($\lesssim 1$ ms) and having final NS masses $\gtrsim 2.5\,{\rm M}_{\odot}$. With a $\beta$ of $0.7$ (or an accretion efficiency of $0.3$), the fastest spinning Galactic BW observed so far can be reproduced (PSR J0952–0607, with a spin period of 1.41 ms; Bassa et al., 2017). However, these short spins correspond to the case where the final NS mass exceeds the maximum NS mass limit. Removing these, the minimum spin period for $\beta=0.7$ is 3.8 ms, which is almost thrice the minimum observed BW spin. Looking at the case of $\beta=0.3$ in Figure 2, most of the observed spins can be reproduced. Comparing the minimum final spin period for $\beta=0.3$ (1.38 ms) and the corresponding final pulsar mass ($2.36\,{\rm M}_{\odot}$) with PSR J0952–0607, we find that our simulations can reproduce well this extreme BW pulsar (its estimated pulsar mass is $2.35^{+0.17}_{-0.17}\,{\rm M}_{\odot}$; Romani et al., 2022).
@@
In Figure 2, there are some binary tracks that ended with final spins below 1 ms, in the sub-MSP regime (all with $M^{i}_{\rm NS}>1.3\,{\rm M}_{\odot}$). The NSs with $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$ gained a maximum of $1.06\,{\rm M}_{\odot}$ and did not spin up to the sub-MSP regime (final spin was 1.37 ms). All the binaries where the final NS mass exceeded the maximum NS mass limit, were spun up to sub-milliseconds. Hence, the mass required to spin up an NS to sub-millisecond range is high enough to collapse it into a BH. Additionally, the sub-millisecond regime is mainly dominated by initially more massive NSs (majority of the $M^{i}_{\rm NS}=2.0$${\rm M}_{\odot}$tracks in Figure 2 appear to be above the maximum NS mass limit) while the initially less massive NSs are not always able to accrete enough matter to enter the sub-millisecond regime. This is because more massive NSs leads to more a stable and longer lasting mass transfer phase (Misra et al., 2020). Comparing the final spin distributions between $\beta=0.3$ and $\beta=0.7$, we can see that $\beta=0.7$ results in slower spinning NS, which would then have lower spin-down luminosities as $L_{\rm pulsar}\propto P^{-3}_{\rm spin}$, and not affecting the orbital evolution as much. Increased wind mass loss from the companion would lead to increased orbital widening during the radio pulsar phase. With $\beta=0.0$, while the entire observed spider parameter space is reproduced, the spin evolution extends down to 1 ms. The value of $\beta$ that is most consistent with the observed minimum pulsar spin is $0.3$. Additionally, $\beta=0.3$ leaves some room for uncertainty in accretion and in our implementation of the X-ray irradiated accretion disk instabilities. With future observations of spiders, any spins approaching 1 ms would signify more conservative nature of RLO at play. Hence, $\beta\lesssim 0.3$ is required to explain the observed spins of spider pulsars.
@@
Figure 3: Evolution of selected binaries from our simulated grids for $\beta=0.3$, $f_{\rm pular}=0.0$, and $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$. The three different kinds of binaries that we discuss in the text are studied here: case A ($P^{i}_{\rm orb}=1.0\rm\,d$), near-$P_{\rm bif}$ ($P^{i}_{\rm orb}=2.4\rm\,d$), and case B or diverging ($P^{i}_{\rm orb}=2.6\rm\,d$). Also shown is the binary that goes down to the orbital period of 19 min (orange curve) for comparison, which is the minimum in the grid with $f_{\rm pulsar}=0.0$. Top left shows the orbital period versus companion mass, where $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. The formation of a fully convective star and the start of the ejection phase are shown as squares. We also compare our tracks to estimated $M_{\rm c,min}$ and observed $P_{\rm orb}$ for BWs (faint black crosses), RBs (faint red crosses), and huntsmen spiders (faint green crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017). The tidarrens, namely, J1311–3430, J1653–0159, and J0636+5128, are highlighted as pink, blue, and black crosses, respectively. Top right, bottom left, and bottom right show the time evolution of the NS spin, companion mass loss rate, and NS mass, respectively, for the same binaries. The time axis is zoomed in to focus on the spin-up phase.
@@ 3.2 Orbital evolution and surface hydrogen abundance
The top left panel in Figure 1 shows the evolutionary tracks with no pulsar wind irradiation ($f_{\rm pulsar}=0.0$). The color bar corresponds to the surface H abundance of the companion, which can be compared to optical spectroscopic observations of spiders. We mark in Figure 1, spiders where H lines have not been detected in their optical spectra (filled triangles) and those where H lines are detected (filled circles). Two of the most compact BWs, PSR J1311–3430 (Pletsch et al., 2012; Romani et al., 2012) and PSR J1653–0159 (Kong et al., 2014; Romani et al., 2014), with negligible surface H abundance, are easily reproduced by our evolutionary tracks. These sources also form a sub-category of BWs, called tidarrens, with $P_{\rm orb}\lesssim 2$ hr and $M_{\rm c}\lesssim 0.015\,{\rm M}_{\odot}$ (Romani et al., 2016). They are reproduced regardless of the value of $f_{\rm pulsar}$. Their evolutionary tracks also pass through the ultra-compact X-ray binary region (UCXB), with orbital periods less than one hour, with the minimum orbital period being 30.8 min. Most UCXBs are also considered to be H-deficient due to their small orbits (Nelson et al., 1986; Nelemans et al., 2006), similar to our binary evolution tracks for tidarrens. Hence, tidarrens (the most compact BWs with the lightest companions) passed through a UCXB phase in their previous evolution. Additionally, we also see these systems reproduced in Figure 13 (with $\beta=0.7$), since not a lot of accretion is needed to form these sources.
@@
We can understand their evolution better in Figure 3 (top left panel) where we show tracks that pass through the tidarren region (solid gray line, with an $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=$2.4 d) and compare these with other tracks with initially narrower and wider orbits. All the tracks shown in the figure have an $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$. We also show a binary that goes well into the UCXB region during its evolution, to reach the overall minimum orbital period (19 min) in the grid slice (with an $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=$2.6 d). With increasing orbital periods, the binaries go from starting RLO with the companion on the MS (at $P_{\rm orb}=0.5$ d, also called case A RLO), to starting RLO when the companion has exhausted core-H (also called case B RLO, at $P_{\rm orb}=1.1$ d). We also see this delay in RLO in bottom right in Figure 3, showing the time evolution of the mass-loss rates of the companions. This transition occurs at the bifurcation period ($P_{\rm bif}$), because companions in wider orbits have more time to evolve and move out of the MS. In the presented example, $P_{\rm bif}$ is between $P_{\rm orb}=0.93$ d (with $P^{i}_{\rm orb}=2.4$ d and $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$) and $P_{\rm orb}=1.1$ d (with $P^{i}_{\rm orb}=2.6$ d and $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$). We term systems just below $P_{\rm bif}$ as near-$P_{\rm bif}$ binaries; similar studies have shown that the formation of UCXBs from LMXBs happens just below $P_{\rm bif}$ (Podsiadlowski et al., 2002, 2003; He et al., 2019; Gossage et al., 2023).
@@
For binary evolution governed only by conservative mass transfer, the orbit would expand as long as the mass is being transferred from the less massive to the more massive star, while it would contract if the companion star is more massive. With the involvement of magnetic braking, this evolution is not as simple. In NS binaries with MS companions having masses around $1\,{\rm M}_{\odot}$, orbital angular momentum evolution is dominated by magnetic braking, which contracts the orbits. Since case A RLO proceeds on a nuclear timescale during the evolution of the stellar MS, any mass loss via winds is not fast enough to dominate over magnetic braking. We can see this nuclear timescale RLO phase in Figure 3 (bottom left panel). In compact binaries with orbital periods $\lesssim 1\rm\,d$, angular momentum losses due to gravitational wave radiation are also significant and further contract the orbit. Once the companion in case A binaries decreases to about 0.2 to 0.3 ${\rm M}_{\odot}$, it develops a fully convective structure and magnetic braking stops operating (Rappaport et al., 1983; Spruit & Ritter, 1983; Pylyser & Savonije, 1989; Podsiadlowski et al., 2002). This is most prominent for the binary in Figure 3 (top left panel) with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=1$ d; there is a decrease in the orbit without any change in mass in the companion mass. The companion detaches from its Roche lobe as the main process driving RLO (the first RLO phase for this binary) has stopped operating, until it fills its Roche lobe again due to gravitational wave radiation continuing to contract the orbit (which is the drop in orbital period seen in the figure). The gravitational wave radiation sustains this second RLO phase and the orbit eventually expands. Correspondingly, we also see this break in the mass-transfer rate in the same figure after which the mass-transfer rate increases again for the second RLO phase. The overall effect of this RLO phase is seen in the top and bottom left panels of Figure 3, case A NSs undergo accretion much more ($\sim 0.85\,{\rm M}_{\odot}$) and have an earlier spin-up than other binaries.
For case B RLO (corresponding to $M^{i}_{\rm c}=1.5$${\rm M}_{\odot}$and $P^{i}_{\rm orb}=2.6$ d in Figure 3), strong stellar winds take away stellar angular momentum. Since angular momentum lost due to stellar winds expands the orbits and the effect of magnetic braking is weaker in comparison, we get diverging orbits. The companion has almost an order of magnitude higher mass-loss rate than case A. The binary is at the very end of the companion MS at the onset of RLO (central H abundance less than 0.1); hence, there is a small contraction in the orbit before it expands during the giant phase. The resulting binaries from these diverging systems are NS/He-WDs in wide orbits. Since this is a shorter-lived phase than case A, the NS does not accrete as much for the same initial donor mass ($\sim 0.4\,{\rm M}_{\odot}$).